OpenStudy (anonymous):

Help Me Please!!!!! F is the midpoint of AB and E is the midpoint of AC If FE=x+5 and BC=3x+2, what is FE? a.)10 b.)11 c.)12 d.)13

OpenStudy (anonymous):



OpenStudy (anonymous):

Please Help

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

@jdoe0001 do you think you'd be able to help me??

OpenStudy (jdoe0001):

http://www.regentsprep.org/regents/math/geometry/gp10/midlinel.htm any ideas?

OpenStudy (anonymous):

it still hard to do, and i looked at that and i don't understand that.... :(

OpenStudy (anonymous):

I just don't understand it at all, and I have looked at alot of examples but can never get it sadly :(

OpenStudy (jdoe0001):

hmm well using the side-splitter theorem hmm lemme check

OpenStudy (jdoe0001):

notice that the midsegment in that link DE is 1/2 of the longer AB

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

 and you have those values thus we konw that FE = x+5 and that BC = 3x+2 thus \(\bf FE=x+5\qquad BC=3x+2\qquad FE=\cfrac{1}{2}BC\to \cfrac{BC}{2}\quad thus \\ \quad \\ x+5=\cfrac{3x+2}{2}\) solve for "x"

OpenStudy (anonymous):

shoot okay so you would end up with x=2 am i right??

OpenStudy (anonymous):

i don't think i did that correct....

OpenStudy (jdoe0001):

well.. is not 2

OpenStudy (anonymous):

okay so i did do that wrong..

OpenStudy (anonymous):

so what is x???